OpenStudy (anonymous):

SOMEONE PLEASE HELP ME QUICK I ONLY HAVE 20 MINTUES TO ANSWER THIS AND ID BE SO GRATEFUL IF SOMEONE ANSWERED There are 30 homes in Neighborhood A. Each year, the number of homes increases by 20%. Just down the road, Neighborhood B has 45 homes. Each year, 3 new homes are built in Neighborhood B.

OpenStudy (anonymous):

Part A: Write functions to represent the number of homes in Neighborhood A and Neighborhood B throughout the years. (4 points)

OpenStudy (anonymous):

Part B: How many homes does Neighborhood A have after 5 years? How many does Neighborhood B have after the same number of years? (2 points)

OpenStudy (anonymous):

Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically. (4 points)

OpenStudy (anonymous):

@Kash_TheSmartGuy can you help please?

OpenStudy (kash_thesmartguy):

Part A: \[f(x)=30*0.2x\]and\[f(x)=45+3x\] Part B: \[f(x)=30+0.2(5)\]\[f(x)=45+3(5)\]For Part B solve them. Part C: No idea. Ask someone else.

OpenStudy (anonymous):

thank you! one last thing can you tell me how to solve part B?

OpenStudy (anonymous):

@kash_thesmartguy

OpenStudy (anonymous):

ill try part C

OpenStudy (anonymous):

wow you're very helpful! haha

OpenStudy (kash_thesmartguy):

You get it, now? Just use formulas

OpenStudy (kash_thesmartguy):

@sparkles456

OpenStudy (anonymous):

@Kash_TheSmartGuy yes because of mrskipped8thgrade

OpenStudy (kash_thesmartguy):

gj @MrSkipped8thGrade

OpenStudy (anonymous):

thank you!!! @MrSkipped8thGrade & no just to justify your answer mathematically

OpenStudy (anonymous):

which you did

OpenStudy (anonymous):

20% of 30 is 6 (for neighborhood A) so every year adds 6 houses

OpenStudy (anonymous):

wait i messed up a few times in there let me backtrack

OpenStudy (anonymous):

1 A=36 B=48 2 A=42 B=51 3 A=48 B=54 4 A=54 B=57 5 A=60 B=60 Answer is actually 5 years exacly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for part B neighborhood A has 60 as well as neighborhood B (after 5 years)

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

for part A neighborhood A f(x)=30+6x neighborhood B f(x)=45+3x where x = number of years

OpenStudy (anonymous):

the first one wouldn't be f(x) = 30 x 0.2x ?

OpenStudy (anonymous):

plug in the formula for after 3 years

OpenStudy (anonymous):

ok i did I'm going to submit now

OpenStudy (anonymous):

f(x) = 30 x 0.2(3) f(x) = 30 x .6 f(x) = 18 this shows a decrease in houses. obviously not it.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem, anytime. :)

OpenStudy (anonymous):

did you get it answered on time? or are you turning an assignment in late? or did you say you only had twenty mins to attract answers?

OpenStudy (anonymous):

@sparkles456 ???????

OpenStudy (anonymous):

i turned it in late.

OpenStudy (anonymous):

@MrSkipped8thGrade

OpenStudy (anonymous):

that sucks oh well

OpenStudy (anonymous):

let me know if you ever have other questions and I will try to help you. :)

OpenStudy (anonymous):

yeah but you were really helpful because i was having a lot of trouble with that lesson

OpenStudy (anonymous):

ok :)